OpenStudy (anonymous):

v^2=v0^2+2ad Solve for a. 0 is the subscript

OpenStudy (anonymous):

vt^2=V0^2+2ad a=vt^2-V0^2/2d

OpenStudy (anonymous):

where does that 2 come from? in both equations?